A Girl is walking along a 45 meter cliff overlooking the River when she kicks a

ball off the ledge with a horizontal velocity of 9.3 m/s. How far away will

the ball land? What will the balls velocity be when it hits the ground?

4. If the girl is 1.97 meters tall, what was the angle of impact the ball had on

the top of her head? What was the balls final velocity when it struck her? (Assume the

ball had an original horizontal velocity of 9.3 m/s)

Please step by step I am confused

I mostly just need help with the second part
1
Expert's answer
2020-11-17T13:03:05-0500

1)

"x=v_xt=9.3\sqrt{\frac{2(45)}{9.8}}=28\ m\\v_y=\sqrt{2(9.8)(45)}=29.7\frac{m}{s}\\v=\sqrt{9.3^2+29.7^2}=31.1\frac{m}{s}""\alpha=\arctan{\frac{9.3}{29.7}}=17.4\degree"